# Benchmarking 3D Human Pose Estimation Models under Occlusions

### v2

@@ 3.2 Benchmark Dataset
To evaluate how each model generalizes to in-the-wild scenarios, we benchmark them on the BlendMimic3D [22] dataset. Since all selected models were trained on the Human3.6M [21] dataset, evaluating them on that same dataset would not assess their robustness to distribution shifts or unseen occlusions. BlendMimic3D provides a more realistic test set, with occlusions caused by object interactions and different viewpoints.
@@
We use subject S2 from BlendMimic3D, who performs 14 actions involving various occludes placed in the scene. To ensure compatibility across models and avoid retraining, we manually adapted the BlendMimic3D skeleton format in Blender to closely match the keypoint structure and conventions of Human3.6M, as can be seen in Figure 2. The converted 2D and 3D pose annotations, along with occlusion labels, will be made publicly available.
Figure 2: Comparison of 2D human pose representations across different formats. (Left) Human3.6M (H36M) format over its original image. (Middle) BlendMimic3D (BM3D) format and its adapted-to-H36M format and (Right) its respective poses.
@@
Figure 3 shows the percentage of occlusion per keypoint across all frames, with an average of $33.93\%$ of keypoints occluded per frame. Notably, RHip and LHip are the most frequently occluded keypoints. However, in the Human3.6M format Hip, RHip and LHip are collinear (as illustrated in Figure 2). This geometric relationship allwos one to infer the third point if two are known. Our results confirm this: despite being consistently occluded, the hip keypoints are still well predicted by 3D HPE models. Figure 3 presents the distribution of occluded keypoints per frame. Notably, there are no frames where all keypoints are visible or completely occluded, the majority of frames have between 3 and 7 occluded keypoints.
@@ 3.3 Occlusion Simulation Protocol
To evaluate the robustness of 3D HPE models under occlusions, we designed two noise injection experiments using the BlendMimic3D dataset. In both cases, we simulate the effect of occlusion by corrupting 2D input poses with Gaussian noise, following the behavior of real 2D detectors under occluded conditions.
@@ 3.3.1 Protocol 1: Adding Noise to Occluded Keypoints
In order to understand how well each model can infer 3D poses from corrupted 2D input data, and thus assess the models’ practical usability in different occlusion levels, we tracked their performances across different noise levels. We applied zero-mean Gaussian noise with varying levels of standard deviation to occluded 2D keypoints, using the per-keypoint occlusion annotations available in BlendMimic3D. This approach allows us to systematically simulate uncertainty in 2D detections without introducing bias or directional shifts.
Let $\mathbf{x}_{ij}\in\mathbb{R}^{2}$ denote the 2D position of keypoint $j$ in frame $i$, and let $\mathcal{O}_{ij}\in\{0,1\}$ be the corresponding occlusion indicator, where $\mathcal{O}_{ij}=1$ denotes that the keypoint is occluded. We define the noise level parameter $\sigma\in\{0,0.001,0.005,0.01,0.03,0.05\}$, which is expressed as a fraction of the average image resolution. The corresponding standard deviation in pixel units is given by
@@
We inject noise into each occluded keypoint by sampling a 2D perturbation vector from an isotropic Gaussian distribution:
@@
This process perturbs only the occluded keypoints in the 2D pose sequence, injecting Gaussian noise independently to the $x$ and $y$ coordinates of each occluded keypoint. We tested six noise levels, $\sigma=\{0,0.001,0.005,0.01,0.03,0.05\}$, applied over 10 independent runs per level for statistical robustness. The $\sigma=0$ case corresponds to ground-truth input. Figure 4 illustrates qualitative examples across noise levels. In our experiments we show that all models are highly sensitive to increased variance, even with zero mean, with notable degradation in 3D pose estimation accuracy.
@@ 3.3.2 Protocol 2: Per-Keypoint Occlusion Analysis
In the second evaluation, we analyze the sensitivity of each model to the occlusion of individual keypoints. The goal is to quantify the relative importance of each keypoint in accurate 3D pose reconstruction. For this, we selected the first 243 frames of each sequence and, over a 100-frame window (frame 50–149), for each of the 17 keypoints, injected zero-mean Gaussian noise into that keypoint alone, keeping all the others unaltered. This was repeated over five independent runs per keypoint to account for variability. By isolating occlusion to a single keypoint at a time, this test allows us to assess which keypoints impact 3D pose accuracy the most. The results help identify which body parts are easier or more difficult for each model to predict.
The standard deviation used for the Gaussian noise was chosen based on findings from the first evaluation and informed by real detector statistics, which are discussed later in the results section. The specific noise level is $\sigma=0.03$ and was introduced after analyzing the baseline sensitivity to ensure that it realistically reflects the magnitude of occlusion-induced errors observed in practice.
@@ 3.4 Evaluation Metrics
Our 3D human pose estimation evaluation is based on the Mean Per-keypoint Positional Error (MPJPE), which measures the average Euclidean distance (in millimeters) between the predicted and ground-truth 3D join positions. MPJPE is represented by the equation
@@
  • Visible MPJPE: The average error computed over all keypoints labeled as visible across all frames. This isolates the model’s performance on clearly detected keypoints.
  • Occluded MPJPE: The average error computed over all keypoints labeled as occluded across all frames. This directly reflects the model’s ability to estimate uncertain keypoints.
All errors are scaled to millimeters and averaged over all actions and test runs for statistical robustness. When applicable, we report standard deviation as a confidence measure across runs.
@@ 4 Results and Discussion
4.1 Occlusion Severity
To assess how 3D HPE models handle increasing occlusion errors, we evaluated their performance using Protocol 1 (Section 3.3.1), where zero-mean Gaussian noise is applied to keypoints labeled as occluded in the BlendMimic3D annotations. We tested six levels of noise, $\sigma$, across ten independent runs. Table 3 reports MPJPE results (in millimeters) for each model.
Table 3: Evaluation of 3D HPE performance for state-of-the-art models under varying levels of simulated occlusion noise on the BlendMimic3D dataset. Results are reported as MPJPE (mm). The best results for each noise level are shown in bold, and the second-best are underlined. DTF-GT corresponds to the original configuration proposed in [20], where occluded keypoints are treated as missing (i.e., zero 2D coordinates) prior to interpolation. No missing keypoints, with simulated noise for the occluded keypoints. No interpolation, considering the ground-truth occlusion labels.
Noise Level
Models Year $\sigma=0$ $\sigma=0.001$ $\sigma=0.005$ $\sigma=0.01$ $\sigma=0.03$ $\sigma=0.05$
VideoPose3D [12] 2019 $73.94\pm 4.85$ $73.93\pm 4.69$ $74.09\pm 4.70$ $74.77$±$\pm 4.74$ $91.11$±$\pm 4.90$ $123.12\pm 7.99$
DiffuPose [17] 2023 $83.45\pm 6.99$ $83.66\pm 6.72$ $88.48\pm 6.69$ $101.20\pm 6.99$ $168.71\pm 12.68$ $223.32\pm 16.59$
FinePose [18] 2024 $69.75$±$\pm 8.87$ $69.78$±$\pm 8.65$ $70.78$±$\pm 9.67$ $77.34\pm 11.12$ $133.05\pm 18.84$ $183.86\pm 23.95$
DTF-GT [20] 2024 $80.83\pm 3.69$
DTF [20] 2024 $80.83\pm 3.69$ $93.39\pm 3.92$ $93.37\pm 3.96$ $93.42\pm 4.01$ $94.14\pm 4.29$ $95.79$±$\pm 4.89$
Despite the noise applied to occluded keypoints being zero-mean, Table 3 reveals that all models exhibit clear sensitivity to increasing standard deviation. This indicates that the evaluated models are vulnerable to noise provided by 2D detectors when a keypoint is occluded.
The performance trends observed in Table 3 can be largely explained by the architectural designs of each model. VideoPose3D, a convolution-based baseline, maintains stable accuracy at low-to-moderate noise levels but lacks explicit mechanisms to handle severe occlusions, leading to a sharper decline under high noise.
P-STMO, despite its transformer-based design, shows limited robustness, likely due to its reliance on masked modeling strategies that may not generalize to real occlusion patterns. MixSTE, which alternates spatial and temporal transformer blocks, performs best in clean settings but degrades quickly under noise, suggesting its structured modeling may be sensitive to keypoint corruption. PoseFormerV2, in contrast, demonstrates strong resilience, especially at moderate noise levels, likely due to its use of frequency-domain features that help mitigate temporal inconsistencies.
D3DP, a diffusion-based method, maintains strong performance under moderate noise, highlighting the advantage of probabilistic modeling, though its accuracy decreases at higher noise levels. DiffuPose, while also diffusion-based, consistently underperforms, potentially due to fragile inference dynamics and limited noise conditioning. FinePose leverages fine-grained prompt conditioning, incorporating action and kinematic-aware guidance, and delivers competitive performance. However, it still lacks explicit mechanisms for occlusion robustness.
In conclusion, models like, MixSTE, D3DP, and PoseFormerV2 perform well under light to moderate occlusion noise, while DTF stands out under severe conditions. Yet none manage to maintain consistent performance across the full occlusion spectrum. This underscores a critical gap in the field: current state-of-the-art models struggle to balance accuracy on clean inputs with robustness to occlusion. Addressing this trade-off should be a central goal for future 3D human pose estimation research.
Although diffusion-based models are often expected to be robust to noisy inputs, this assumption does not fully hold here. All three tested methods, D3DP, DiffuPose, and FinePose, condition their denoising process directly on the 2D input pose, which guides every step. As a result, noise in occluded keypoints significantly degrades their predictions, revealing a consistent weakness in current diffusion designs when faced with realistic occlusions. Moreover, these models are computationally expensive, requiring iterative sampling during inference, which further limits their practicality in real-time.
[…]
As a complementary analysis to the results presented in Table 3, Figure 5 shows the normalized MPJPE, computed as the difference between each noise level and the baseline at $\sigma=0$, for three error types: overall pose error (average MPJPE), occluded keypoint error, and visible keypoint error. This visualization highlights how model performance degrades under increasing occlusion noise and reveals differences in sensitivity to visible versus occluded keypoints.
Figure 5: Normalized MPJPE across increasing noise levels. Each curve represents the MPJPE for overall, occluded, and visible keypoints, normalized by the respective model’s baseline performance at $\sigma=0$. (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF.
Across all models, errors on occluded keypoints increase more steeply than on visible ones, revealing a common lack of explicit occlusion modeling. In contrast, DTF exhibits almost flat curves, with both visible and occluded MPJPE remaining stable across noise levels. MixSTE and FinePose show the greatest sensitivity to noise, particularly in occluded keypoints, where errors rise sharply related to visible keypoints.
DiffuPose suffers the most degradation across all error types, indicating poor robustness in its inference dynamics. D3DP maintains relative stability for visible keypoints but is notably affected by noise in occluded keypoints.
VideoPose3D, P-STMO and PoseFromerV2 remain relatively stable under mild noise ($\sigma<0.01$). Among them, P-STMO degrades faster, while VideoPose3D shows a more abrupt error increase as noise intensifies. PoseFormerV2 demonstrates the highest resilience, likely due to its frequency-domain temporal encoding, which helps mitigate the impact of occluded keypoint noise and supports more balanced performance.
4.1.1 DTF Performance
To accurately assess DTF’s behavior under realistic occlusion scenarios, we introduced three distinct test configurations detailed in Table 4. These correspond directly to the DTF results reported in Table 3. The original DTF model (DTF-GT) expects exactly 16 keypoints to be randomly occluded per frame. In their code, these occluded keypoints are set to zero (i.e., treated as missing), and their positions are then estimated through interpolation, guided by visibility-based confidence scores derived from temporal context. However, their assumption of having only one visible keypoint at a time and random occlusion patterns is completely unrealistic. In our evaluation using BlendMimic3D, which provides naturally occurring occlusion annotations, keypoints can remain occluded for longer durations than DTF expects, as can be seen in Figure 6. Such prolonged occlusions render DTF’s interpolation strategy less effective, indicating that its training approach is inadequate for real-world occlusions.
Occl. Labels Zero 2D Coord. Interp. Lifting
DTF-GT
DTF
DTF
Figure 6: Analysis of the duration of occlusions per keypoint over the actions, from the ground-truth occlusion labels of BlendMimic3D subject S2, (a) the average duration and (b) maximum duration.
To bridge this gap between simulated and realistic occlusion, we evaluated two modified variants of the original DTF model: (1) DTF, which uses full 2D poses (including noisy occluded keypoints) and applies interpolation using the ground-truth occlusion labels, and (2) DTF†‡, which removes interpolation and directly inputs the noisy poses with occlusion labels. Table 4 summarizes the differences among these variants in terms of input preparation and processing.
[…]
On the other hand, configurations using raw noisy poses with occlusion labels (DTF, DTF†‡) yield more stable predictions, even as noise increases. Despite being designed for missing keypoints, DTF demonstrates robustness to real noisy poses, performing best at $\sigma=0.05$. This resilience stems from its dual-view generation and fusion architecture, along with a confidence-aware interpolation mechanism that enables the model to down-weight unreliable keypoints and leverage temporal consistency effectively.
4.1.2 T3D-CNN Performance
Unlike DTF, T3D-CNN was designed to work with 2D keypoints masked as missing (set to zero), relying heavily on an internal occlusion mask provided as input. To evaluate its robustness, we trained T3D-CNN under three different masking strategies: no masking (None) , 8 randomly occluded keypoints (Rand 8) and 16 randomly occluded keypoints (Rand 16). Our goal was to assess the model’s ability to generalize across varying numbers of missing keypoints.
Table 5 presents the performance of each model under multiple test conditions: 2D ground-truth input from Human3.6M with no missing keypoints, with 8 or 16 randomly masked keypoints, and ground-truth 2D input from BlendMimic3D dataset, which includes realistic occlusions but lacks the original occlusion mask format expected by T3D-CNN.
[…]
Test Configuration
H36M
No miss Rand 8 Rand 16 BM3D
Model None 358.14
Rand 8 153.50 56.79 419.15 314.65
Rand 16 74.42
The results indicate that T3D-CNN struggles to generalize beyond the specific masking condition it was trained on. Performance degrades when tested on unseen masking configurations or on BM3D, where the number and duration of occluded keypoints vary naturally based on the scene and motion. Notably, the model trained with 8 randomly occluded keypoints (Rand 8) performs better on BM3D than the versions trained with no masking or 16 occlusions. This suggests that the configuration of Rand 8 is more reasonable and better aligned with the range of occlusion patterns found in BM3D, as shown in Figure 3. Because T3D-CNN cannot process complete or noisy 2D inputs, and lacks a mechanism to exploit real occlusion annotations, it was excluded from the noise-level evaluation presented in Table 3.
4.1.3 Occlusion Noise Analysis
@@
After evaluating model robustness across a range of Gaussian noise levels (Protocol 1), we wanted to relate these simulations in real-world conditions. Specifically, we aimed to understand whether the selected noise levels realistically reflect the types and magnitudes of errors produced by standard 2D detectors under occlusion. Therefore, we conducted an error analysis of two commonly used 2D detectors, CPN [60] and Detectron2 [71], on 1000x1002 images from subject S2 in BlendMimic3D. The detectors output 2D poses in COCO format [23], which we aligned semantically with the Human3.6M-converted ground-truth, excluding non-aligned keypoints such as head, neck, spine, eyes and ears. Table 6 presents the MPJPE and standard deviation for both visible and occluded keypoints. Occlusions were further categorized into two groups: self-occlusions, where a subject occludes themselves, and external occlusions, caused by objects or out-of-frame positioning.

### v3

@@ 3.2 Benchmark Dataset
To assess generalization to in-the-wild scenarios, we benchmark models on the BlendMimic3D [24] dataset. Since all selected models were trained on Human3.6M [15] dataset, evaluating them on that same dataset would not assess their robustness to distribution shifts or unseen occlusions. BlendMimic3D provides a more realistic test set, with occlusions caused by object interactions and different viewpoints.
@@
We use subject S2 from BlendMimic3D, who performs 14 actions with various objects placed in the scene. To ensure compatibility across models and avoid retraining, we manually adapted the BlendMimic3D skeleton format in Blender to closely match the keypoint structure and conventions of Human3.6M, as can be seen in Figure 2. The converted 2D and 3D pose annotations, along with occlusion labels, will be made publicly available.
Figure 2: Comparison of 2D human pose representations across different formats. (Left) Human3.6M (H36M) format over the original image. (Middle-Left) BlendMimic3D (BM3D) format and its version adapted to H36M. (Middle-Right) The corresponding 2D poses. (Right) The corresponding 3D poses.
@@
Figure 3 shows the percentage of occlusion per keypoint across all frames, with an average of $33.93\%$ of keypoints occluded per frame. Notably, RHip and LHip are the most frequently occluded keypoints. In the Human3.6M format, Hip, RHip, and LHip are collinear (as illustrated in Figure 2), which makes it difficult in Blender to accurately segment the body volume of each keypoint individually. As a result, these keypoints are often labeled as occluded. However, our results show that, despite frequent occlusion, hip keypoints are still well predicted by 3D HPE models. Figure 3 presents the distribution of the number of occluded keypoints per frame. Notably, there are no frames where all keypoints are visible or completely occluded, the majority of frames have 3 to 7 occluded keypoints.
@@ 3.3 Occlusion Simulation Protocol
To evaluate the robustness of 3D HPE models under occlusions, we designed two noise injection experiments using the BlendMimic3D dataset. In both cases, we simulate the effect of occlusion by corrupting 2D input poses with Gaussian noise, following the behavior of real 2D detectors.
@@ 3.3.1 Protocol 1: Adding Noise to Occluded Keypoints
In order to understand how well each model can infer 3D poses from corrupted 2D input data, and thus assess the models’ practical usability in different occlusion levels, we tracked their performances across different noise levels. We applied zero-mean Gaussian noise with varying levels of standard deviation to occluded 2D keypoints, using the per-keypoint occlusion annotations available in BlendMimic3D. This approach allows us to systematically simulate errors in 2D detections without introducing bias.
Let $\mathbf{x}_{ij}\in\mathbb{R}^{2}$ denote the 2D position of keypoint $j$ in frame $i$, and let $\mathcal{O}_{ij}\in\{0,1\}$ be the corresponding occlusion indicator, where $\mathcal{O}_{ij}=1$ denotes that the keypoint is occluded. We define the noise level parameter $\sigma\in\{0.001,0.005,0.01,0.03,0.05\}$, which is expressed as a fraction of the average image resolution. The corresponding standard deviation in pixel units is given by
@@
We add noise into each occluded keypoint by sampling a 2D perturbation vector from an isotropic Gaussian distribution:
@@
This process perturbs only the occluded keypoints in the 2D pose sequence, injecting Gaussian noise independently to the $x$ and $y$ coordinates of each occluded keypoint. We tested five noise levels, $\sigma=\{0.001,0.005,0.01,0.03,0.05\}$, applied over 10 independent runs per level for statistical robustness. For comparison, we also included the case of 2D ground-truth pose input. Figure 4 illustrates qualitative examples across noise levels.
@@ 3.3.2 Protocol 2: Per-Keypoint Occlusion Analysis
In the second evaluation, we analyze the sensitivity of each model to the occlusion of individual keypoints. The goal is to quantify the relative importance of each keypoint in accurate 3D pose estimation. For this, we selected the first 243 frames of each sequence and, over a 100-frame window (frame 50–149), for each of the 17 keypoints, injected zero-mean Gaussian noise into that keypoint alone, keeping all the others unchanged. This was repeated over five independent runs per keypoint to account for variability. By isolating occlusion to a single keypoint at a time, this test allows us to assess which keypoints impact 3D pose accuracy the most. The results help identify which body parts are easier or more difficult for each model to predict.
The standard deviation used for the Gaussian noise was chosen based on findings from the first evaluation and informed by real detector statistics, which are discussed later in the results section. This ensures that the simulated noise realistically reflects the typical errors caused by occlusions in real-world scenarios.
@@ 3.4 Evaluation Metrics
Our evaluation is based on the Mean Per-keypoint Positional Error (MPJPE), which measures the average Euclidean distance (in millimeters) between the predicted and ground-truth 3D keypoint positions. MPJPE is represented by the equation
@@
  • Visible MPJPE: The average error computed over all keypoints labeled as visible across all frames. This isolates the model’s performance on visible keypoints.
  • Occluded MPJPE: The average error computed over all keypoints labeled as occluded across all frames. This directly reflects the model’s ability to estimate occluded keypoints.
All errors are reported in millimeters and averaged over all actions and test runs for statistical robustness. When applicable, we report standard deviation as a confidence measure across runs.
@@ 4 Results and Discussion
4.1 2D Detector Error Analysis
@@
To better understand the behavior of 2D pose detectors under occlusion, we conducted an analysis focusing on the types and magnitudes of detection errors in realistic scenarios. To this end, we analyzed two commonly used 2D detectors, CPN [7] and Detectron2 [50], on 1000x1002 images from subject S2 in BlendMimic3D. The detectors output 2D poses in COCO format [23], which we aligned semantically with the Human3.6M-converted ground-truth, excluding non-aligned keypoints such as head, neck, spine, eyes and ears. Table 3 presents the MPJPE and standard deviation for both visible and occluded keypoints. Occlusions were further categorized into two groups: self-occlusions, where a subject occludes themselves, and external occlusions, caused by objects or out-of-frame positioning.
@@
Additionally, as shown in Table 3, the standard deviation of 2D detection errors for visible keypoints is generally below $1\text{\,}\%$ of the image resolution, while occluded keypoints show a higher spread, reaching up to $2.37\text{\,}\%$. Consequently, we interpret values of $\sigma=\{0.001,0.005,0.01\}$ as representative of noise from visible keypoints, and values $\sigma=\{0.03,0.05\}$ as simulating the broader uncertainty caused by occlusions.
4.2 Protocol 1: Adding Noise to Occluded Keypoints
To assess how 3D HPE models handle increasing occlusion errors, we evaluated their performance using Protocol 1 (Section 3.3.1), where zero-mean Gaussian noise is applied to keypoints labeled as occluded in the BlendMimic3D annotations. The selected models span four categories: convolutional-based [32], transformer-based [38, 57, 59], diffusion-based [39, 9, 53], and occlusion-aware [11, 12] methods.
To better structure our analysis, we divide the discussion into two parts. First, we analyze models without explicit occlusion handling mechanisms (convolutional, transformer-based, and diffusion-based). Then, we study occlusion-aware models, designed to handle occluded keypoints through architectural or training strategies.
4.2.1 Models Without Occlusion Awareness
Figure 5: Normalized MPJPE across increasing noise levels. Each curve represents the MPJPE for overall, occluded, and visible keypoints, normalized by the respective model’s baseline performance using 2D ground-truth poses. (a) VideoPose3D, (b) P-STMO, (c) MixSTE, (d) PoseFormerv2, (e) D3DP, (f) DiffuPose, (g) FinePose, and (h) DTF. These plots align with the quantitative results presented in Table 4.
Table 4 reports the MPJPE results for all evaluated models except T3D-CNN [12]. We exclude T3D-CNN from this comparison because the model was originally designed to treat occluded keypoints as missing, setting their 2D coordinates to zero. This behavior is incompatible with our benchmark setting, where occluded keypoints retain noisy values, and leads to invalid or misleading results. For this reason, we evaluate T3D-CNN separately in the following subsection, where we analyze the model with its masking-based occlusion mechanism.
Table 4: Evaluation of 3D HPE performance for state-of-the-art models under varying levels of simulated occlusion noise on the BlendMimic3D dataset. Results are reported as MPJPE ($m$). The best results for each noise level are shown in bold, and the second-best are underlined. DTF-GT corresponds to the original configuration proposed in [11], where occluded keypoints are treated as missing (i.e., zero 2D coordinates) prior to interpolation. No missing keypoints, with simulated noise for the occluded keypoints. No interpolation, considering the ground-truth occlusion labels.
Noise Level
Models Year GT $\sigma=0.001$ $\sigma=0.005$ $\sigma=0.01$ $\sigma=0.03$ $\sigma=0.05$
@@ 4.2.1 Models Without Occlusion Awareness
DTF-GT [11] 2024 $80.83\pm 3.69$
[…]
Despite Gaussian noise being zero-mean, the models included in Table 4 exhibit clear sensitivity to increasing standard deviation. This indicates that the evaluated models are vulnerable to noise provided by 2D detectors when a keypoint is occluded.
Focusing on models without occlusion-awareness [32, 57, 38, 59, 39, 9, 53], the performance trends observed in Table 4 can be largely explained by the architectural designs of each model. VideoPose3D, a convolution-based baseline, maintains stable accuracy at low-to-moderate noise levels but lacks explicit mechanisms to handle severe occlusions, leading to a sharper decline under high noise.
P-STMO, despite its transformer-based design, shows limited robustness, likely due to its reliance on masked modeling strategies that may not generalize to real occlusion patterns. MixSTE, which alternates spatial and temporal transformer blocks, performs best in clean settings but degrades quickly under noise, suggesting its structured modeling may be sensitive to keypoint corruption. PoseFormerV2, in contrast, demonstrates strong resilience, especially at moderate noise levels, likely due to its frequency-domain temporal encoding, which helps mitigate the impact of occluded keypoint noise and supports more balanced performance.
D3DP, a diffusion-based method, maintains strong performance under moderate noise, highlighting the advantage of probabilistic modeling. It uses MixSTE as a denoiser backbone, conditioned only on the 2D input pose, which guides every step. As a result, noise in occluded keypoints significantly degrades their predictions. DiffuPose, while also diffusion-based, consistently underperforms. It uses a Graph Convolutional Network (GCN) denoiser, likewise conditioned solely on the 2D input pose, which may limit its robustness to noisy inputs. FinePose achieves competitive performance through fine-grained prompt conditioning, incorporating action and kinematic-aware guidance. Nonetheless, its denoiser is also conditioned on 2D poses. Thus, the lack of explicit mechanisms for handling occlusions in these diffusion-based models results in degraded performance at higher noise levels ($\sigma$). Moreover, these models are computationally expensive, requiring iterative sampling during inference, which further limits their practicality in real-time.
Overall, the results in Table 4 suggest that PoseFormerV2 is particularly resilient to moderate occlusion noise, outperforming both older models like VideoPose3D and newer diffusion-based approaches such as FinePose.
As a complementary analysis to the results presented in Table 4, Figure 5 shows the normalized MPJPE, computed as the difference between each noise level and the baseline (i.e., using ground truth 2D poses), for three error types: overall pose error (average MPJPE), occluded keypoint error, and visible keypoint error. This visualization highlights how model performance degrades under increasing occlusion noise and reveals differences in sensitivity to visible versus occluded keypoints.
Across all models without occlusion-awareness, errors on occluded keypoints increase more steeply than on visible ones, revealing the common lack of explicit occlusion modeling. MixSTE and FinePose show the greatest sensitivity to noise, particularly in occluded keypoints, where errors rise sharply relative to visible keypoints.
DiffuPose suffers the most degradation across all error types, indicating poor inference robustness. D3DP maintains relative stability for visible keypoints but is notably affected by noise in occluded keypoints.
VideoPose3D, P-STMO and PoseFormerV2 remain relatively stable under mild noise ($\sigma<0.01$). Among them, P-STMO degrades faster, while VideoPose3D shows a more abrupt error increase as noise intensifies, and PoseFormerV2 demonstrates the highest resilience.
Collectively, these models are sensitive to occluded keypoints. This sensitivity is especially evident in diffusion-based architectures since their denoisers are conditioned directly on the 2D input at every iteration, any corruption introduced by occlusions propagates through the entire sampling process. Future diffusion-based methods should explore alternative conditioning mechanisms, uncertainty-aware denoisers, or hybrid formulations. In contrast, PoseFormerV2’s resilience highlights the potential of frequency-domain representations, which suppress high-frequency jitter produced by occlusions. Nonetheless, all evaluated models show that existing training strategies are insufficient to prepare models for real-world scenarios. Occlusion-aware learning, explicit visibility modeling, or architectures capable of reasoning over uncertain inputs could improve robustness.
4.2.2 Occlusion-Aware Models
Unlike the models discussed previously, DTF [11] and T3D-CNN [12] incorporate explicit mechanisms to address occlusions. These models are evaluated separately to better understand how occlusion-aware components.
DTF Performance: DTF handles occlusion through preprocessing and architecture design. During preprocessing, occluded keypoints have their 2D coordinates set to zero (i.e., treated as missing) and are temporally interpolated using the nearest visible instance of the same keypoint within a window spanning the previous 40 frames and the next 40 frames (81-frame window). If no such instance exists, the keypoint remains zero. Each keypoint receives also a confidence score reflecting its visibility or interpolation uncertainty. These (x, y, confidence) inputs are then fed to the model. During inference, the confidence values inform the transformer about the reliability of each joint. The architecture itself is a dual-view transformer that generates two intermediate 3D hypotheses, refines them with self-attention, and fuses them with cross-view attention to produce the final 3D pose.
During training the authors randomly occlude 16 keypoints per frame, and their positions are then estimated through interpolation. However, their assumption of having only one visible keypoint at a time and random occlusion patterns is unrealistic. In our evaluation using BlendMimic3D, which provides naturally occurring occlusion annotations, keypoints can remain occluded for longer durations than DTF expects, as can be seen in Figure 6. Such prolonged occlusions render DTF’s interpolation strategy less effective, indicating that its training approach is inadequate for real-world occlusions.
Figure 6: Analysis of occlusion duration per keypoint across actions. Based on ground-truth occlusion labels from BlendMimic3D (subject S2), this figure reports the average and maximum occlusion durations for each keypoint.
To bridge this gap between simulated and realistic occlusion, we evaluated two modified variants of the original DTF model (DTF-GT): (1) DTF, which uses full 2D poses (including noisy occluded keypoints) and applies interpolation using the ground-truth occlusion labels, and (2) DTF†‡, which omits interpolation entirely and instead takes the noisy poses with occlusion labels as input. Table 5 summarizes the differences among these variants in terms of input preparation and processing. These correspond directly to the DTF results reported in Table 4.
Occl. Labels Zero 2D Coord. Interp. Lifting
DTF-GT
DTF
DTF
In Table 4, the DTF-GT configuration fails when occlusion is present ($\sigma=\{0.001,0.005,0.01,0.03,0.05\}$). This occurs because the model treats occluded keypoints as missing (i.e., zeroed 2D coordinates) and attempts to recover it through interpolation within a $\pm 40$-frame window. However, as shown in Figure 6, BlendMimic3D contains realistic occlusion that often last longer than 81 frames, meaning many keypoints have no visible instance within the search window. Therefore these joints remain zeroed, a scenario that almost never occurs in the original DTF training setup. Since the model is not trained to handle prolonged missing joints, it collapses and outputs the same high error ($301.91\pm$34.60\text{\,}\mathrm{m}\mathrm{m}$\$) across all noise levels. These results show that while the original interpolation strategy may be effective for synthetic occlusion patterns, it struggles with the duration and structure of realistic occlusions in BlendMimic3D, suggesting the need to extend the temporal window or replacing interpolation with a learned occlusion-aware reconstruction of the 2D pose.
[…]
T3D-CNN Performance: Unlike DTF, T3D-CNN is explicitly designed to operate on 2D keypoints where occluded joints are masked to zero, using a binary occlusion-guidance input to indicate which joints are missing. As a temporal dilated CNN, it relies on this mask to inpaint the missing joints and lift the sequence to 3D under a specific occlusion setting. To evaluate its robustness, we trained T3D-CNN under three different masking strategies: no masking (None), 8 randomly occluded keypoints (Rand 8) and 16 randomly occluded keypoints per frame (Rand 16). Our goal was to assess the model’s ability to generalize across varying numbers of missing keypoints.
Table 6 presents the performance of each model under multiple test conditions: 2D ground-truth input from Human3.6M with no missing keypoints, with 4, 8, 12 or 16 randomly masked keypoints, and 2D input from BlendMimic3D dataset, where keypoints are zeroed, providing realistic occlusions, though not in the fixed random-mask format expected by T3D-CNN.
Table 6: Evaluation of T3D-CNN under varying conditions. All inputs use 2D ground-truth keypoints in both Human3.6M (H36M) and BlendMimic3D (BM3D). >1000 indicates that the error exceeds $1000\text{\,}\mathrm{m}\mathrm{m}$. Rand 4, Rand 8, Rand 12 and Rand 16 refer to random occlusion of 4, 8, 12 or 16 keypoints, respectively, applied during training or testing as indicated. The best results for each test configuration are shown in bold.
Test Configuration
H36M
No miss Rand 4 Rand 8 Rand 12 Rand 16 BM3D
Model None 41.20 274.23 352.97 403.38 457.04 529.61
Rand 8 153.50 100.78 56.79 256.89 419.15 314.65
Rand 16 >1000 >1000 >1000 907.93 74.42 >1000
For the None model, the two occlusion-guidance input channels of T3D-CNN were removed,
as the network was trained using only fully visible keypoints.
The results indicate that T3D-CNN struggles to generalize beyond the specific masking condition it was trained on. Because the model is always trained with a fixed number of missing keypoints (e.g., exactly 16 per frame), its filters and batch-norm statistics become specialized to that sparsity level. When evaluated on unseen masking configurations, such as a different number of occlusions or the naturally varying patterns in BM3D, the occlusion-mask distribution shifts, placing the model in a regime it was never trained for, leading to large performance drops. A more robust design would require training with a mixture of occlusion patterns so the model learns to handle a broader range of scenarios.
Notably, as shown in Table 6, the model trained with 8 randomly occluded keypoints (Rand 8) maintains more stable performance across a wider range of missing keypoints when evaluated on the H36M dataset. This behavior is consistent with the fact that it performs better on BM3D than the versions trained with no masking or 16 occlusions. This also suggests that the configuration of Rand 8 is more reasonable and better aligned with the range of occlusion patterns found in BM3D, as shown in Figure 3.
Because T3D-CNN cannot process complete or noisy 2D inputs, and lacks a mechanism to exploit real occlusion annotations, it was excluded from the noise-level evaluation presented in Table 4.
4.3 Protocol 2: Per-Keypoint Occlusion Analysis